Conference On L^{A}TEX 2ε

2000/02/21(Mon)

• 10:00–10:50 AAA From L^{A}TEX2.09 to L^{A}TEX 2ε.

• 11:00–11:50 BBB Using graphics with L^{A}TEX 2ε. EPS, BMP and JPEG.

• 14:00–14:50 CCC Using Plain TEX Commands—Why I use them.

• 15:00–15:50 DDD Why XML—What is it?

• 16:10–17:00 EEE L^{A}TEX 2ε on Internet.

2000/02/22(Tue)

• 9:30–10:20 FFF

From  To  .

• 10:40–12:10 GGG(Invited Lecturer) Virtual Fonts Technology, Part I.

• 13:00–13:50 HHH Using .

• 14:00–14:50 III Using TrueType Fonts. Times New Roman, Arial and Courier New.

• 15:00–15:50 HHH How to Prepare Presentation Manuscript on DVIOUT.

• 16:00–17:30 Short Communications.

2000/02/23(Wed)

• 9:30–11:00 GGG Virtual Fonts Technology, Part II.

• 11:10–12:00 JJJ How to Write New Macro Package.

• 13:00–13:50 KKK Structure of L^{A}TEX Class Files.

• 14:00–14:50 LLL New Features of L^{A}TEX 2ε*.

• 15:00–16:30 MMM How to Prepare Beautiful Manuscript on L^{A}TEX.

• 16:40–17:30 NNN Future L^{A}TEX.